OpenStudy (anonymous):

can someone please give me the quadratic form of x/x-1 - the square root of x/x-1 raised to the 6th power - 40=0

OpenStudy (anonymous):

Can someone help me check my answers (6 questions on the Law of Sines)? a few moments agoReport Abuse Share Tweet This Question is Open 47 mangnagelytis Best Response 0 a few moments agoReport Abuse 2 attachments untitled.png untitled2.png 100 satellite73 Best Response 0 which one? a few moments agoReport Abuse 70 Eyad Best Response 0 @satellite73 Please when you finish this one ,May u check my question ? Ty :) a few moments agoReport Abuse 100 satellite73 Best Response 0 @Eyad sorry i don't know a spot of physics a few moments agoReport Abuse 70 Eyad Best Response 0 @satellite73 : oh ,its k then Ty for your time *Statics* a few moments agoReport Abuse 100 satellite73 Best Response 0 don't know that either a few moments agoReport Abuse 100 satellite73 Best Response 0 @mangnagelytis your C is right 31.47 a few moments agoReport Abuse 47 mangnagelytis Best Response 0 Sorry, I'm back, it's all six. a few moments agoReport Abuse 99 radar Best Response 0 2nd link, question 6. It doesn't say which weather station, you have solved for one distance, I assume you have to solve the other distance. You do have the data to solve it.

OpenStudy (anonymous):

report is done

OpenStudy (mertsj):

\[\frac{x}{x-1}-(\sqrt{\frac{x}{x-1}})^6-40=0\]

OpenStudy (mertsj):

Is that it?

OpenStudy (anonymous):

Yes

OpenStudy (mertsj):

\[(\sqrt{\frac{x}{x-1}})^6=(\frac{x}{x-1})^3\]

OpenStudy (mertsj):

So if you have posted it accurately I can't see a quadratic form for it.

OpenStudy (anonymous):

Oh, well that's ok. I can ask my teacher tomorrow about it, but thanks anyway.

OpenStudy (mertsj):

@ fullmetal777 Are you sure it's a square root or is it a third root?

OpenStudy (anonymous):

It said square root, but I think my classmate found the answer

OpenStudy (mertsj):

What is it?